# Active Fine-Tuning of Generalist Policies

### v1

@@ 4.2 Practical Algorithms
Nevertheless, we can leverage the approximation of neural networks as a linear functions over an embedding space $\bm{\pi}(s,c;\theta)=\bm{\beta}^{\top}\phi_{\theta}(s,c)$, where both weights $\bm{\beta}$ and embeddings $\phi_{\theta}(\cdot)$ exist in a $p$-dimensional latent space (Lee et al., 2019; Khan et al., 2019). This technique does not violate any of the practical constraints listed above, and allows us to adapt the machinery introduced in GP settings. While several embedding strategies exist Jacot et al. (2018); Devlin et al. (2019); Holzmüller et al. (2023), we adopt loss gradient embeddings (Ash et al., 2020a). Assuming the prior $\bm{\beta}\sim\mathcal{N}(0,I)$, the policy $\bm{\pi}(s,c;\theta)$ can be modeled by a Gaussian Process with kernel $k_{\theta}((s,c),(s^{\prime},c^{\prime}))=\langle\phi_{\theta}(s,c),\phi_{\theta}(s^{\prime},c^{\prime})\rangle$. When coupled with this approximation, the conditional entropy objective in Equation 2 can be reformulated:
$c_{n}=\operatorname*{arg\,max}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}),\;\tau^{\prime}\sim\bm{\tau}(c^{\prime})\\ c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\end{subarray}}\sum_{t=0}^{H-1}k_{\theta}((s_{t},c),X)[k_{\theta}(X,X)+\sigma_{\epsilon}^{2}I]^{-1}k_{\theta}(X,(s_{t},c)),$ (4)
where $X$ is vector of states and tasks in $(c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1})$. As the collected dataset grows, the conditioning on previous trajectories $\tau_{1:n-1}$ can instead be addressed by fine-tuning the network’s parameters $\theta$ (e.g., through conventional gradient descent), resulting in updates in the embedding function $\phi_{\theta}$.
Batch selection
Standard practice in deep active learning prescribes the collection of batches of samples at each round (Gal et al., 2017; Sener & Savarese, 2018; Ducoffe & Precioso, 2018). This necessity is partially addressed by the fact that, in our setting, a single demonstration contains several samples for training the policy (as many as the length $H$ of the demonstration). Nevertheless, we can further leverage the GP approximation to select multiple demonstrations at each round. In fact, a simple recursive greedy selection can provide a constant factor approximation on the information gain objective (Krause & Golovin, 2014; Hübotter et al., 2024). In practice, when selecting $M$ demonstrations at the $n$-th round, we can select the $m$-th demonstration out of $M$ by applying the criterion in Equation 1, while making sure to condition entropy estimates on both the $m-1$ demonstrations already selected in this round, and the $M\cdot(n-1)$ demonstrations previously collected.
@@
A fundamental issue with neural function approximation under shifts to the training distribution is known as forgetting (McCloskey & Cohen, 1989; French, 1999), as any further optimization may catastrophically perturb pre-trained parameters. Common strategies for its mitigation often involve rehearsal (Atkinson et al., 2021; Verwimp et al., 2021) or regularization (Kirkpatrick et al., 2017). Unfortunately, the former is not feasible in this setting due to lack of access to pre-training data, and the latter was not found to be empirically effective (see Appendix G). Scale and a diverse pre-training dataset can also mitigate forgetting (Dyer et al., 2022), but neither can be controlled during fine-tuning. Thus, we opt for a parameter isolating solution (Rusu et al., 2016; Yu et al., 2020), in which the task-space is partitioned (e.g. uniformly), and a copy of the fine-tuning parameters is stored and trained for each partition, thus avoiding negative interference. Inference can then be performed by selecting the parameter set through a nearest-neighbor lookup in task-space. This solution can be easily applied for limited task sets, and can be scaled to large task spaces through discretization schemes. While parameter isolation prevents constructive interference across tasks, we found it to bring a net benefit during fine-tuning.
By combining the approximations required by AMF-GP with three additional design choices (namely kernel approximation, batch selection and parameter isolation), we obtain a method for active multi-task fine-tuning of policies parameterized via neural networks. We refer to this practical instantiation as AMF-NN.
Figure 2: 2D integrator. Starting from the origin, each task involves reaching a given point on a circle, as shown in differently colored trajectories.
@@ 5 Experiments
The experiment section is designed to evaluate active multi-task fine-tuning and provide an empirical answer to several questions. We thus reserve a section to each of them.
@@ 5.1 When is AMF beneficial?
Uniform pre-training (12/12 tasks)
  AMF-GP    Uniform
Skewed pre-training (6/12 tasks)
Figure 3: Experiments in GP settings for a 2D integrator (see Figure 2). AMF-GP selects tasks that minimize the policy’s posterior entropy and improves the agent’s returns faster than uniform task sampling. In the middle, the improvement in final return over the baseline is greater when the pre-training distribution is skewed and includes fewer tasks. We report return and entropy curves for non-skewed and skewed pre-training (left and right, respectively). We report means and 90% simple bootstrap confidence intervals over 10 random seeds ; dots and crosses mark corresponding measurements.
When none of the assumptions listed in Section 4.1 is violated, AMF is guaranteed to converge to the optimal policy. We furthermore investigate whether AMF also results in faster empirically faster convergence with respect to naive approaches to data collection. To do so, we compare AMF to uniform i.i.d. sampling from the set of tasks $\mathcal{C}$. First, we consider a classic 2D integrator as a benchmark environment (see Figure 2). The agent is a pointmass initialized in the origin, and can directly control its 2D velocity, which is integrated over the past trajectory to return the current state. We can define an infinite task space, in which each task consists of reaching a point on a circle centered on the origin, and the agent is rewarded with the negative Euclidean distance to it. The evaluation distribution $\mu_{c}$ assigns equal probability to 12 points in different directions. The initial state distribution is deterministic, dynamics are both deterministic and smooth, while the expert policy is smooth and corrupted with i.i.d. Gaussian noise. We model the policy as a Gaussian Process with a RBF kernel, and we condition it on a pre-training dataset of $12$ noisy demonstrations. We then collect $50$ additional demonstrations by running both AMF-GP and uniform sampling.
We first consider a perfectly uniform pre-training regime, in which each evaluation task is demonstrated exactly once (Figure 3, left). As the policy’s entropy is minimized, AMF-GP increases the policy’s returns at a higher rate compared to uniform sampling of demonstrations. We then extend this evaluation to several pre-training distributions (Figure 3, middle), and compare the final performance of the two methods as the pre-training budget is allocated to a decreasing number of tasks. As the pre-training distribution becomes more skewed (e.g., when only 6/12 tasks are demonstrated in Figure 3, left), we observe that the performance gap between uniform task sampling and AMF-GP grows larger. This is to be expected, as in this case the information gain from the next demonstration heavily depends on the queried task, and taking the $\operatorname*{arg\,max}$ of the criterion in Equation 1 is significantly better than choosing a random task. Intuitively, in this case, uniform sampling of tasks fails to reliably provide demonstrations for tasks that were observed less often during pre-training.
@@
0510Rounds0.40.50.60.7Multi-task success rateKitchen, uniform
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
0510Rounds0.40.50.60.70.8Multi-task success rateMetaworld, uniform
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
@@
0510Rounds0.200.250.300.350.40Multi-task success rateKitchen, uniform
0510Rounds0.150.200.250.300.350.40Multi-task success rateKitchen, skewed
0510Rounds0.50.60.70.80.9Multi-task success rateMetaworld, uniform
0510Rounds0.40.60.8Multi-task success rateMetaworld, skewed
  AMF-NN ($\sigma^{2}=10^{-4}$)   AMF-NN ($\sigma^{2}=10^{-3}$)    AMF-NN ($\sigma^{2}=10^{-2}$)
  Uniform    Uniform, no parameter isolation
Figure 4: AMF with neural policies in Frankakitchen (top) and Metaworld (bottom). Experiments are repeated for state and RGB inputs (left and right). We evaluate both uniform and skewed pre-training distribution. AMF-NN is overall desirable, and highly beneficial for skewed pre-training distributions. We report means and 90% simple bootstrap confidence intervals over 10 seeds.
@@ 5.2 Can AMF scale to high-dimensional tasks?
In both environments, we evaluate AMF-NN when learning from state measurements, as well as from raw pixels. In the first case, the policy is simply parameterized through a MLP, while in the second the MLP receives the embedding of a pre-trained visual encoder (Nair et al., 2022b). The policy is pre-trained on $\approx 15$ total demonstrations, which we allocate either uniformly across all tasks, or only on half of them, reproducing the uniform and skewed regimes from the previous experiments. Afterwards, we apply AMF-NN for $10$ iterations, collecting $2$ demonstrations at each iteration.
Figure 4 reports average multi-task success rates at each iteration compared to a random uniform task selection scheme. For the baseline, we additionally report performance without parameter isolation, highlighting performance degradation due to negative gradient interference across tasks. We observe that AMF’s performance is on par or better than a uniform task sampling scheme when the prior was trained uniformly on all tasks. However, as reported in the previous section, AMF is very beneficial when the pre-training dataset does not uniformly cover the evaluation tasks. These trends are consistent across both environments, and both modalities.
Finally, Figure 4 also highlights the effect of the only hyperparameter introduced by AMF-NN, namely the noise parameter $\sigma^{2}$ in the GP approximation. While selecting an excessively high noise level results in slightly greedier behavior and premature convergence, we overall observe that the method is not particularly sensitive to this hyperparameter, performing fairly reliably across 2 orders of magnitude.
5.3 Is batch-wise task selection important for AMF?
In the GP setting, policy fine-tuning and conditioning for batch-wise selection correspond. Thus, it is not necessary to collect batches of demonstrations. This is not the case for AMF-NN. On one hand, collecting batches of training data is computationally beneficial since training can be parallelized effectively (Sener & Savarese, 2018). On the other hand, each query already returns several samples in our setting, which could make batch selection unnecessary. We thus set out to empirically validate which batch size is desirable for active data collection in our setting.
4.04.55.05.56.06.5Area under success rate curveKitchen, uniform
4.04.55.05.56.06.5Area under success rate curveKitchen, skewed
$B\!=1$ $B\!=2$ $B\!=4$ Uniform
Uniform, no parameter isolation
Figure 5: Influence of batch size $B$ over area under success rate curve with 20 demonstrations.
0510Rounds0.550.600.650.700.75Multi-task success rateKitchen, uniform
0510Rounds0.40.50.60.70.8Multi-task success rateKitchen, skewed
  Last-layer gradient
  Last-layer    Dropout
@@ 5.3 Is batch-wise task selection important for AMF?
While keeping the total budget fixed to 20 demonstrations, we evaluate AMF-NN with batch sizes $B$ spanning from 1 to 4, and report the area under the success rate curve in Figure 5. For convenience, we report the same metric for the uniform selection baselines from Figure 4, for which the data selection strategy does not depend on the batch size. We find that larger batch sizes are not necessarily desirable.
@@ 5.4 How do uncertainty estimates for AMF compare?
As entropy estimation is at the core of AMF-NN, we additionally compare the adopted GP approximation with loss-gradient embeddings to other approaches from the literature. In praticular, we also consider an alternative GP approximation using last-layer embeddings (Holzmüller et al., 2023), as well as test-time Dropout (Loquercio et al., 2020). For the latter, each batch is simply filled with demonstrations from the task maximizing prior entropy, that is $\operatorname*{arg\,max}_{c\in\mathcal{C}}\mathop{\mathbb{E}}_{\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1}),\;(s_{0},\dots)\sim\bm{\tau}(c)}\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid\tau_{1:n-1})$. Both of these schemes are in practice desirable, as they do not require access to action labels. However, we observe that these two schemes are prone to early convergence to suboptimal task choices, or are less effective in driving task selection. Hence, as shown in Figure 6, multi-task performance is in general lower, suggesting that the entropy estimation technique is crucial to AMF-NN.
@@ 5.5 Can AMF be applied to off-the-shelf models?
  AMF-NN    Uniform
  Uniform, no parameter isolation
024Rounds0.400.450.500.550.600.65Multi-task success rateWidowX
As AMF-NN has minimal requirements (essentially, access to a differentiable pre-trained prior is sufficient), it should be widely applicable. In this section, we investigate scaling our evaluation to recently published open-source generalist policies. For this purpose, we chose Octo (Octo Model Team et al., 2024). This model relies on a transformer backbone for integrating multimodal information (in the form of state sensors, camera images and text or RGB task descriptions), and uses a diffusion-based policy head for action prediction (Chi et al., 2023). For computational reasons, we will focus on fine-tuning the action head alone. Octo is pre-trained on a large-scale real-world robotic dataset (Collaboration, 2023), and is thus designed for inference on physical hardware. Nonetheless, a recently proposed evaluation suite enables simulated evaluations that statistically correlate with real-world results (Li et al., 2024). We thus collect rollouts from a pre-trained Octo on the WidowX tasks, and filter them to only include successes, akin to self-distillation schemes Bousmalis et al. (2024). On availability of such self-supervised demonstrations, we then apply AMF-NN for $4$ iterations, providing $4$ demonstrations in each round. The results are reported in Figure 7.
As all evaluation tasks are largely demonstrated in the pre-training dataset (Collaboration, 2023), we find that AMF-NN does not improve significantly upon uniform task collection, confirming the trend we observed for uniform pre-training distributions in Figure 4. Nonetheless, we observe that it constitutes an effective method for data selection, and can be applied as a drop-in replacement for fine-tuning of off-the-shelf models. [^3] This evaluation also reports an interesting trend, that is a vast reduction in catastrophic forgetting, to the point that parameter isolation is not necessary. This anecdotal evidence can be seen as an instance of a general trend of mitigated catastrophic forgetting in large models (Dyer et al., 2022).

### v3

@@ 4.2 Practical Algorithms
Nevertheless, we can leverage the approximation of neural networks as a linear functions over an embedding space $\bm{\pi}(s,c;\theta)=\bm{\beta}^{\top}\phi_{\theta}(s,c)$, where both weights $\bm{\beta}$ and embeddings $\phi_{\theta}(\cdot)$ exist in a $p$-dimensional latent space (Lee et al., 2019; Khan et al., 2019). This technique does not violate any of the practical constraints listed above, and allows us to adapt the machinery introduced in GP settings. While several embedding strategies exist (Jacot et al., 2018; Devlin et al., 2019; Holzmüller et al., 2023), we adopt loss gradient embeddings (Ash et al., 2020). Assuming the prior $\bm{\beta}\sim\mathcal{N}(0,I)$, the policy $\bm{\pi}(s,c;\theta)$ can be modeled by a Gaussian Process with kernel $k_{\theta}((s,c),(s^{\prime},c^{\prime}))=\langle\phi_{\theta}(s,c),\phi_{\theta}(s^{\prime},c^{\prime})\rangle$. When coupled with this approximation, the conditional entropy objective in Equation 2 can be reformulated:
$c_{n}=\operatorname*{arg\,min}_{c^{\prime}\in\mathcal{C}}\mathop{\mathbb{E}}_{\begin{subarray}{c}c\sim\mu_{c},\;(s_{0},\dots)\sim\bm{\tau}(c)\\ \tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})\\ \tau^{\prime}\sim\bm{\tau}(c^{\prime})\end{subarray}}\sum_{t=0}^{H-1}K(s_{t},c,X),$ (4)
where $K(s_{t},\!c,\!X)\!:=\!k_{\theta}(x_{t},\!X)[k_{\theta}(X,\!X)\!+\!\sigma_{\epsilon}^{2}I]^{-1}k_{\theta}(X,\!x_{t})$, $x_{t}=(s_{t},c)$, and $X$ is the vector of states and tasks in $(c^{\prime},\tau^{\prime},c_{1:n-1},\tau_{1:n-1})$. As the collected dataset grows, the conditioning on previous trajectories $\tau_{1:n-1}$ can instead be addressed by fine-tuning the network’s parameters $\theta$ (e.g., through conventional gradient descent), resulting in updates in the embedding function $\phi_{\theta}$.
No mismatch (12/12 tasks demonstrated)
  AMF-GP    Uniform
Mismatch (6/12 tasks demonstrated)
Figure 2: Experiments in GP settings for a 2D integrator (see Figure 3). AMF-GP selects tasks that minimize the policy’s posterior entropy and improves the agent’s returns faster than uniform task sampling. In the middle, the improvement in final return over the baseline is greater when the pre-training distribution does not perfectly match the evaluation distribution $\mu_{\mathcal{C}}$ and only demonstrates fewer tasks. We report return and entropy curves for non-mismatched and mismatched pre-training (left and right, respectively). We plot means and 90% simple bootstrap confidence intervals over 10 random seeds ; dots and crosses mark corresponding measurements.
@@
Catastrophic forgetting is an inherent challenge to neural function approximation under shifts to the training distribution, as optimization over new data is non-local, and may undo learning progress (McCloskey & Cohen, 1989; French, 1999). This issue is critical in our setting, as actively guiding the fine-tuning distribution will necessarily accentuate the distribution shift from pre-training. Common strategies for its mitigation often involve rehearsal (Atkinson et al., 2021; Verwimp et al., 2021) or regularization (Kirkpatrick et al., 2017). Unfortunately, the former is not possible in this setting due to lack of access to pre-training data, and the latter was not found to be empirically effective (see Appendix H). Scale and a diverse pre-training dataset can also mitigate forgetting (Ramasesh et al., 2022), but neither can be controlled during fine-tuning.
We thus propose a practical algorithmic solution to alleviate catastrophic forgetting when fine-tuning multi-task policies. Intuitively, we would like the policy to retain the skills it mastered during pre-training, while focusing on novel tasks during fine-tuning. Inspired by previous works in behavioral priors (Bagatella et al., 2022) and offline RL (Kumar et al., 2020), we propose retain a copy of the pre-trained policy, which we refer to as prior ($\bm{\pi}_{p}$). For a given state $s\in\mathcal{S}$ and task $c\in\mathcal{C}$, we can then linearly combine actions sampled from the fine-tuned policy $\bm{\pi}$ with those sampled from the prior: $a=\alpha(c)\hat{a}+(1-\alpha(c))\bar{a}$, where $\hat{a}\sim\bm{\pi}(\cdot|s,c),\bar{a}\sim\bm{\pi}_{\text{p}}(\cdot|s,c)$. Crucially, $\alpha(c)\in[0,1]$ is a task-dependent weight trained through gradient descend on an proxy behavior cloning loss, with an additional conservative penalty that encourages closeness to 00. As a result, actions will drift towards the fine-tuned policy’s output as soon as it robustly outperforms pre-training performance on a given task. Tasks which are not improved during fine-tuning will rely instead on samples from the prior, and will not be forgotten. We refer to this technique as Adaptive Prior, and present a detailed description and a comparison to continual learning techniques in Appendix H.
By combining the approximations required by AMF-GP with the described solutions for estimating entropy and preventing catastrophic forgetting, we obtain a method for active multi-task fine-tuning of policies parameterized via neural networks, which we refer to as AMF-NN.
@@ 5 Experiments
The experiment section is designed to evaluate active multi-task fine-tuning and provide an empirical answer to several questions. We thus reserve a section to each of them. Additional evaluations are reported in Appendix D, E and F.
@@ 5.1 When is AMF beneficial?
When none of the assumptions listed in Section 4.1 is violated, AMF is guaranteed to converge to the optimal policy. We furthermore investigate whether AMF also results in faster empirically faster convergence with respect to naive approaches to data collection. To do so, we compare AMF to uniform i.i.d. sampling from the set of tasks $\mathcal{C}$. First, we consider a classic 2D integrator as a benchmark environment (see Figure 3). The agent is a pointmass initialized in the origin, and can directly control its 2D velocity, which is integrated over the past trajectory to return the current state. We can define a continuous task space, in which each task consists of reaching a point on a circle centered on the origin, and the agent is rewarded with the negative Euclidean distance to it. The evaluation distribution $\mu_{c}$ assigns equal probability to 12 points in different directions. The initial state distribution is deterministic, dynamics are both deterministic and smooth, while the expert policy is smooth and corrupted with i.i.d. Gaussian noise. We model the policy as a Gaussian Process with a RBF kernel, and we condition it on a pre-training dataset of $12$ noisy demonstrations. We then collect $40$ additional demonstrations by running both AMF-GP and uniform sampling.
Figure 3: 2D integrator. Starting from the origin, each task involves reaching a given point on a circle, as shown by differently colored trajectories.
@@
01020Rounds0.20.40.60.8Multi-task success rateKitchen, mismatch
01020Rounds0.600.650.700.750.800.85Multi-task success rateKitchen, no mismatch
01020Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, mismatch
01020Rounds0.30.40.50.6Multi-task success rateMetaworld, no mismatch
@@
0510Rounds0.10.20.30.40.5Multi-task success rateKitchen, mismatch
0510Rounds0.200.250.300.350.400.450.50Multi-task success rateKitchen, no mismatch
0510Rounds0.30.40.50.60.7Multi-task success rateMetaworld, mismatch
0510Rounds0.30.40.50.6Multi-task success rateMetaworld, no mismatch
  AMF-NN   Uniform, with adaptive prior   Uniform
Figure 4: AMF with neural policies in Frankakitchen (top) and Metaworld (bottom). Experiments are repeated for state and RGB inputs (left and right). We evaluate both mismatched and non-mismatched settings. AMF-NN is overall desirable, and highly beneficial for mismatched pre-training distributions. We report means and 90% simple bootstrap confidence intervals over 10 seeds.
As a sanity check, we first consider a perfectly uniform pre-training regime, in which each evaluation task is demonstrated exactly once. The pre-training task distribution $\mu_{\mathcal{D}}\in\Delta(\mathcal{C})$ thus perfectly matches the evaluation task distribution $\mu_{\mathcal{C}}$ (Figure 2, left). As it actively minimizes the policy’s entropy, AMF-GP increases the policy’s returns slightly faster when compared to uniform sampling of demonstrations. We then extend this evaluation to more realistic pre-training distributions, characterised by a mismatch with respect to the evaluation distribution: $\mu_{\mathcal{C}}\neq\mu_{\mathcal{D}}$ (Figure 2, middle), and compare the final performance of the two methods as the pre-training budget is allocated to a decreasing number of tasks. As the pre-training distribution diverges from $\mu_{\mathcal{C}}$ (e.g., when only 6/12 tasks are demonstrated in Figure 2, right), we observe that the performance gap between uniform task sampling and AMF-GP grows larger. This is to be expected, as in this case the information gain from the next demonstration heavily depends on the queried task, and taking the $\operatorname*{arg\,max}$ of the criterion in Equation 1 is significantly better than choosing a random task. Intuitively, in this case, uniform sampling of tasks fails to reliably provide demonstrations for tasks that were observed less often during pre-training.
@@ 5.2 Can AMF scale to high-dimensional tasks?
In both environments, we evaluate AMF-NN when learning from state measurements, as well as from raw pixels. In the first case, the policy is simply parameterized through a MLP, while in the second the MLP receives the embedding of a pre-trained visual encoder (Nair et al., 2022). The policy is pre-trained on $\approx 15$ total demonstrations, which we allocate either uniformly across all tasks, or only on half of them, reproducing the non-mismatched and mismatched regimes from the previous experiments. Afterwards, when learning from state measurements, we apply AMF-NN for $20$ iterations, collecting one demonstrations at each iteration. To compensate for the increased complexity, in visual settings we instead provide 2 demonstrations per each task selected and only evaluate 10 iterations.
01020Rounds0.30.40.50.60.70.80.9Multi-task success rateKitchen, mismatch
01020Rounds0.600.650.700.750.800.85Multi-task success rateKitchen, no mismatch
  Last-layer gradient     Last-layer     Dropout
@@
05Rounds0.30.40.50.60.7Multi-task success rateRobomimic, mismatch
05Rounds0.30.40.50.60.7Multi-task success rateRobomimic, no mismatch
  AMF-NN    Uniform, with adaptive prior    Uniform
Figure 6: Evaluation on Robomimic. AMF-NN is widely applicable (e.g., to diffusion policies).
Figure 4 reports average multi-task success rates at each iteration compared to a random uniform task selection scheme. For the baseline, we additionally report performance with an adaptive prior, highlighting how this solution prevents performance degradation due to catastrophic forgetting of pre-training demonstrations. As reported in the previous section, AMF is beneficial under distribution mismatch, in which the pre-training dataset does not exactly cover the evaluation distribution $\mu_{\mathcal{C}}$. As a sanity check, we also observe that, when pre-training demonstrates all tasks equally and a uniform task allocation would be very effective, AMF’s performance matches this naive baseline.
These trends are consistent across both environments, and both modalities. For a qualitative analysis of the strategy induced by AMF, we refer to Appendix I and J.
@@ 5.3 How do uncertainty estimates for AMF compare?
As entropy estimation is at the core of AMF-NN, we additionally compare the adopted GP approximation with loss-gradient embeddings to other approaches from the literature. In particular, we also consider an alternative GP approximation using last-layer embeddings (Holzmüller et al., 2023), as well as test-time Dropout (Loquercio et al., 2020). The latter simply selects the task maximizing prior entropy, that is $\operatorname*{arg\,max}_{c\in\mathcal{C}}\mathop{\mathbb{E}}\sum_{t=0}^{H-1}\mathcal{H}(\bm{\pi}(s_{t},c)\mid\tau_{1:n-1})$, with $\tau_{1:n-1}\sim\bm{\tau}(c_{1:n-1})$ and $(s_{0},\dots)\sim\bm{\tau}(c)$. Both of these schemes are in practice desirable, as they do not require access to action labels. However, we observe that these two schemes are less effective in driving task selection. Hence, as shown in Figure 6 (and in Appendix G), multi-task performance is in general lower, suggesting that the entropy estimation technique is important for AMF-NN.
@@ 5.4 Can AMF be applied to off-the-shelf models?
As AMF-NN has minimal requirements (essentially, access to a differentiable pre-trained prior is sufficient), it should be widely applicable. In this section we investigate further scaling our evaluation to more complex tasks, multi-modal demonstrators and modern policy classes. We consider the Robomimic benchmark, which involves four long-horizon, precise manipulation tasks (up to $\approx$700 steps). While experiments in previous sections rely on demonstrations collected by scripted policies or RL agents, expert trajectories are in this case provided by humans; due to the increased scale, we sample 20 demonstrations for the task selected at each iteration. Finally, we fine-tune a larger generative model, namely a diffusion policy (Chi et al., 2023), which remains compatible with AMF-NN.
Despite the change in data source and architectures, the results we observe in Figure 6 are consistent with those reported in previous settings: active fine-tuning and an adaptive prior are overall helpful, especially when the pre-training distribution does not match the evaluation distribution $\mu_{\mathcal{C}}$.